# What To Do (and Not to Do) with Causal Panel Analysis under Parallel Trends:

### v2

@@ 5.  Systematic Assessment
Other issues.
Our reanalysis highlights several additional issues. First, as we have hinted throughout this section, a large number of studies either lack sufficient statistical power or are on the borderline of being underpowered. According to the authors’ original specifications, one (3%) and 11 (29.7%) studies are not statistically significant at the 5% and 1% levels, respectively. As previously noted, 10 studies lose statistical significance at the 5% level once we apply an HTE-robust estimator. The cause is a combination of a loss of power and, in some cases, smaller point estimates. While only a small fraction of studies in our sample (five articles, 13.5%) employ a bootstrap procedure to estimate SEs or CIs, we observe that the widely practiced cluster-robust SEs typically perform adequately because the numbers of clusters/units are usually fairly large.
Second, the presence of missing values is widespread. Although most methodological work presumes balanced panels without missing data, in reality, many empirical studies encounter varying degrees of data missingness. We plot patterns of treatment status as we do in Figure 2(a) for each study in the SM. Based on these plots, we see that in some studies, the pattern of missingness is either nonrandom or extremely prevalent, which calls into question the reliability and validity of the respective empirical findings.
Third, we perform carryover effect tests for all studies that feature treatment reversals. If this test fails, it suggests that the treatment effects persist beyond the treatment periods. Among 22 articles, five reject the null hypothesis of no carryover effects at 5%. LWX (2022) note that the presence of carryover effects for a limited number of periods is less concerning, as researchers can recode treatment to persist for some time after a unit transitions out of treatment. By adopting this guideline, we observe that carryover effects do not substantially alter the findings in most studies. Specifically, estimates are similar in magnitude, and results initially deemed statistically significant continue to hold their statistical significance (Figure A7 in the SM). Nevertheless, we recommend that researchers make it a practice to check for potential carryover effects, considering the low cost of conducting such tests and adjustments.
Many studies also exhibit sensitivity to model specifications. When we incorporate an LDV or ULT into the authors’ original specifications, a large number of studies lose their statistical significance (29% and 41% respectively). [^14] Plümper and Troeger (2019) raise a cautionary note against the blanket use of fixed effects models in panel data analysis. They show that if both the treatment and the outcome variables are highly autoregressive, these models may lead to spurious correlations. While it is true that such models may be vulnerable to additional biases, such as the Nickell bias in short dynamic panels, and the additional parameters reduce efficiency, these results underscore that a significant number of empirical findings in the literature rely heavily on the modeling assumptions. Hence, we recommend researchers carefully assess the robustness of their findings using different model specifications, possibly in combination with HTE-robust estimators. We reserve a more careful and rigorous analysis of this issue for future studies.
Relatedly, some studies that we exclude from our sample employ one-way FE or FE at a level different from that at which treatment is assigned. Many of these findings do not hold when we reanalyze them using a TWFE model. We should clarify that this does not imply that the original results are not credible; rather, it underlines the fact that these models operate under distinct identification assumptions, and there is substantial variation in how much consideration authors give to this point. We notice that some studies do not provide a rationale for their choice to use one-way fixed effects, while others explicitly outline the type of unobserved confounders they intend to control for. In one instance, the authors inaccurately label their specification as a DID design. We emphasize that TWFE and DID are generally not equivalent. This difference becomes even more pronounced when the fixed effects are not assigned at the level of treatment.
Summary.
In Table 4 below, we summarize the main findings of our reanalysis. The first three columns indicate the proportion of studies in each journal, and across all journals, that have potential PTA violations. The numbers represent the proportion of studies where the null hypothesis is rejected at the 5% significance level for the $F$ test and the placebo test and not rejected for the equivalence test. Across all journals, around a quarter of studies reject the null hypothesis using the $F$ test, with slightly fewer rejecting when using the placebo test, indicating potential PTA violations. Note, though, that failure to reject to reject may result from insufficient power. In 47% of the cases, the equivalence test fails to reject the null hypothesis using $|\widehat{ATT}|$ as the threshold, implying that for these studies we cannot confidently state that the pre-treatment residual averages lie within a narrow range defined by the estimated ATT.
Table 4: Summary of Findings
Potential PTA violations Consequence of HTE Insufficient power
(1) (2) (3) (4) (5) (6) (7)
$F$ test Placebo test Equiv. test FEct-TWFE FEct-TWFE TWFE not FEct not
Journal reject null reject null not reject null Ratio $<0.8$ Ratio $<0.5$ reject null reject null
APSR (6) 0.33 0.33 0.50 0.00 0.00 0.00 0.17
AJPS (12) 0.25 0.17 0.42 0.15 0.00 0.00 0.15
JOP (19) 0.19 0.12 0.50 0.32 0.11 0.05 0.42
All (37) 0.24 0.17 0.47 0.21 0.05 0.03 0.29
Note: A null is deemed being rejected if $p<0.05$. “Ratio” refers to the ratio of the FEct estimate to TWFE estimate for the average effect. Four studies with only a single pre-treatment period are not included in the summary statistics in columns (1)–(3).
Columns (4) and (5) show the ratio of the FEct estimator to the TWFE estimator, which proxies the consequences of HTE. Across all journals, in 21% of studies, this ratio is less than 0.8, and in 5% of studies, the ratio is less than 0.5, indicating that while HTE might be a significant concern theoretically, its practical influence on yielding qualitatively different findings may not be as substantial as the existing literature might suggest. Lastly, columns (6) and (7) show the proportion of studies in which the null hypothesis of no effect is not rejected by TWFE and FEct respectively. When employing the TWFE method, this happens in 3% of the studies across all journals; however, when the FEct estimator is utilized, this number increases to 29%, suggesting many studies in our sample are potentially underpowered.
When we synthesize these multiple lines of evidence, we find that 12 studies (or 32% of our sample) provide strong evidence supporting the plausibility of the PTA—reflected by the rejection of the equivalence test—and distinguish the ATT from zero with sufficient statistical power. If we instead consider studies where the $F$ test has a $p$-value greater than 0.05 (acknowledging that the $F$ test is often underpowered), this number increases to 16 studies (42%).
We emphasize that our goal is not to pass judgment on each individual paper, but rather to understand the state of the literature at large. We thus do not attempt to label ambiguous cases that neither suffer from clear PTA violations and weighting problems nor provide strong evidence to the contrary. Similarly, we do not tabulate subjective calls on whether pretrends look “good” or “bad,” despite our advice for researchers to do so themselves. In practice, when readers are attempting to evaluate the strength of evidence for a specific study, we remind readers that credibility is not binary and urge readers to take a holistic view of the evidence.
6.  What To Do and (Not) To Do with Causal Panel Analysis
We conclude by sharing practical advice based on both our findings and our experiences conducting a vast number of replications and reanalyses using observational panel data. Table 5 summarizes these lessons. The first consideration is the research design. Our results echo the advice from Rubin (2008): “Design trumps analysis,” and it is imperative to understand the underlying assignment mechanism. While fixed effects allow us to control for certain unobserved confounders, this comes at the cost of strong assumptions about functional form as well as the dynamics between the treatment and outcome, such as the absence of anticipation, feedback, and time-varying confounders. If a glaring violation is already known to exist at this stage, a new identification strategy is needed.
[…]

### v3

@@ A.1.6 Choosing Between Estimators
A.2 Proof of Remark A.1
First, we note that de Chaisemartin and D’Haultfœuille (2020) define $\hat{\delta}^{M}$ to allow for ‘group’ level fixed effects that may be higher up than the unit level. Let $N_{g,t}$ denote the number of observations in group $g$ at time $t$. We assume a “sharp design,” meaning all units in the same cell $(g,t)$ have the same treatment. Let $N_{d,d^{\prime},t}=\sum_{g:\begin{subarray}{c}D_{g,t}=d,\\ D_{g,t-1}=d^{\prime}\end{subarray}}N_{g,t}$ denote the number of observations with treatment status $d$ in period $t$ and status $d^{\prime}$ in period $t-1$. Let $Y_{.,g,t}=\frac{1}{N_{g,t}}\sum^{N_{g,t}}_{i=1}Y_{i,g,t}$ denote the average outcome (across observations) in group $g$ at time $t$. Define the following quantities:
$DID_{+,t}$ $=\sum_{g:D_{g,t}=1,D_{g,t-1}=0}\frac{N_{g,t}}{N_{1,0,t}}\big{(}Y_{.,g,t}-Y_{.,g,t-1}\big{)}-\sum_{g:D_{g,t}=D_{g,t-1}=0}\frac{N_{g,t}}{N_{0,0,t}}\big{(}Y_{.,g,t}-Y_{.,g,t-1}\big{)}\text{ and}$
$DID_{-,t}$ $=\sum_{g:D_{g,t}=D_{g,t-1}=1}\frac{N_{g,t}}{N_{1,1,t}}\big{(}Y_{.,g,t}-Y_{.,g,t-1}\big{)}-\sum_{g:D_{g,t}=0,D_{g,t-1}=1}\frac{N_{g,t}}{N_{0,1,t}}\big{(}Y_{.,g,t}-Y_{.,g,t-1}\big{)},$
letting $DID_{+,t}=0$ whenever $\min\{N_{1,0,t},N_{0,0,t}\}=0$ and $DID_{-,t}=0$ whenever $\min\{N_{1,1,t},N_{0,1,t}\}=0$. Finally, define
$\hat{\delta}^{M}=\sum^{T}_{t=2}\Bigg{(}\frac{N_{1,0,t}}{N_{S}}DID_{+,t}+\frac{N_{0,1,t}}{N_{S}}DID_{-,t}\Bigg{)},$
where $N_{S}\vcentcolon=|(g,t):t\geq 2,D_{g,t}\neq D_{g,t-1}|$ is the number of switchers.
Now, we consider the case where the group level is the same as the unit level. Note that then $N_{g,t}=1$ always.
We can now write
$DID_{+,t}=\sum_{i:D_{i,t}=1,D_{i,t-1}=0}\frac{1}{N_{1,0,t}}\big{(}Y_{i,t}-Y_{i,t-1}\big{)}-\sum_{i:D_{i,t}=D_{i,t-1}=0}\frac{1}{N_{0,0,t}}\big{(}Y_{i,t}-Y_{i,t-1}\big{)}$
and similarly
$DID_{-,t}=\sum_{i:D_{i,t}=D_{i,t-1}=1}\frac{1}{N_{1,1,t}}\big{(}Y_{i,t}-Y_{i,t-1}\big{)}-\sum_{i:D_{i,t}=0,D_{i,t-1}=1}\frac{1}{N_{0,1,t}}\big{(}Y_{i,t}-Y_{i,t-1}\big{)}$
Now consider $\hat{\delta}^{PM}$ with the choice of $l=1$, which estimates the contemporaneous treatment effect at the moment of joining treatment,
$\hat{\delta}^{PM}_{1,a}=\frac{\sum_{i=1}^{N}\sum_{t=a+1}^{T}\mathbf{1}\left\{|\mathcal{M}_{it}|>0\right\}D_{i,t}(1-D_{i,t-1})\big{(}(Y_{i,t}-Y_{i,t-1})-\sum_{i^{\prime}\in\mathcal{M}_{i,t}}\frac{1}{|\mathcal{M}_{i,t}|}(Y_{i^{\prime},t}-Y_{i^{\prime},t-1})\big{)}}{\sum_{i=1}^{N}\sum_{t=a+1}^{T}\mathbf{1}\left\{|\mathcal{M}_{it}|>0\right\}D_{i,t}(1-D_{i,t-1})}.$
Now further restrict lags used for matching to $a=1$. Then the matched set $\mathcal{M}_{i,t}=\big{\{}i^{\prime}:i^{\prime}\neq i,D_{i,^{\prime}t}=0,D_{i,^{\prime}t-1}=D_{i,t-1}\big{\}}$ is just units that have the same treatment status in the previous period and are in control in the current period. Under the assumption that a stable group exists, the matched set must be nonempty for any “joiner” ($(i,t):J_{i,t}=1$), where $J_{i,t}=\mathbf{1}\left\{D_{i,t}=1\right\}\mathbf{1}\left\{D_{i,t-1}=0\right\}$, and so $\mathbf{1}\left\{|\mathcal{M}_{it}|>0\right\}D_{i,t}(1-{}_{i,t})=J_{i,t}$. Let $N_{J}\vcentcolon=|(g,t):t\geq 2,J_{i,t}=1|$ be the number of joiners.
Now we have,
$\small\hat{\delta}^{PM}_{1,1}=$ $\frac{\sum_{t\geq 2}\sum_{i:J_{i,t}=1}\big{(}(Y_{i,t}-Y_{i,t-1})-\frac{1}{|\mathcal{M}_{i,t}|}\sum_{i^{\prime}\in\mathcal{M}_{i,t}}(Y_{i^{\prime},t}-Y_{i^{\prime},t-1})\big{)}}{N_{J}}$
$=$ $\frac{1}{N_{J}}\sum_{t\geq 2}\sum_{i:J_{i,t}=1}\Big{(}(Y_{i,t}-Y_{i,t-1})-\frac{1}{|\mathcal{M}_{i,t}|}\sum_{i^{\prime}:D_{i^{\prime},t}=D_{i^{\prime},t-1}=0}(Y_{i^{\prime},t}-Y_{i^{\prime},t-1})\Big{)}$
$=$ $\frac{1}{N_{J}}\sum_{t\geq 2}\sum_{i:J_{i,t}=1}\Big{(}(Y_{i,t}-Y_{i,t-1})-\frac{1}{N_{0,0,t}}\sum_{i^{\prime}:D_{i^{\prime},t}=D_{i^{\prime},t-1}=0}(Y_{i^{\prime},t}-Y_{i^{\prime},t-1})\Big{)}$
$=$ $\frac{1}{N_{J}}\sum_{t\geq 2}\bigg{(}\sum_{i:J_{i,t}=1}(Y_{i,t}-Y_{i,t-1})-\frac{N_{1,0,t}}{N_{0,0,t}}\sum_{i^{\prime}:D_{i^{\prime},t}=D_{i^{\prime},t-1}=0}(Y_{i^{\prime},t}-Y_{i^{\prime},t-1})\bigg{)}$
$=$ $\frac{N_{1,0,t}}{N_{S}}DID_{+,t}.$
We can alter the definition of the matched set to target “leavers” $\{(i,t):D_{i,t}=0,D_{i,t-1}=1\}$ to get an estimate for the contemporaneous effect of leaving $\hat{\delta}^{PM-ART}_{1,1}$ and similarly show that $\hat{\delta}^{PM-ART}_{1,1}=\frac{N_{0,1,t}}{N_{L}}DID_{-,t}$, where $N_{L}$ is the number of leavers. Observe that $\hat{\delta}^{M}=\frac{N_{J}}{N_{S}}\hat{\delta}^{PM}_{1,1}+\frac{N_{L}}{N_{S}}\hat{\delta}^{PM-ART}_{1,1}$.
A.3 Sample and Replicability
A.3.1 Sample Selection Criteria
We collect our replication sample from three leading journals in political science, APSR, AJPS, and JOP. We screen all full research articles published in these journals during 2017-2022 using the following four criteria:
  1. 1.
    The paper uses panel data analysis as a critical piece of evidence to support a causal argument. Specifically, either the abstract or the introduction of the paper needs to mention the results from the panel analysis.
  2. 2.
    The paper uses at least one linear model to analyze panel data, such as DID, TWFE, or lagged dependent variable (LDV) models, and the treatment variable has to be binary. In other words, papers that use only discrete outcome models or continuous treatments are excluded. We include this criterion because most of the analytical tools the literature has developed so far are designed for linear models with binary treatments.
  3. 3.
    We exclude papers that use a regression discontinuity design or an instrumental variables (IVs) design, including Bartik IVs, as their primary identification strategy.
  4. 4.
    We exclude papers that do not exploit within-unit variation despite the longitudinal structure of the data. These designs are drastically different from the rest of the panel studies in their estimand, their identification assumptions, and the properties of their estimators and are worth investigating separately.
A.3.2 Replicability
For papers that meet our four screening criteria, we try to find replication materials from public data-sharing platforms, such as the Harvard Dataverse, and the authors’ personal websites. For each paper, we choose one model that we think can best represent the paper’s central claim. Specifically, we sequentially go through the following two criteria: (1) the authors claim that it is the preferred model; and (2) the model uses the most complete dataset (i.e., with the least missing values). Using data and code from the replication materials, we are able to successfully replicate the main results of 37 of 52 papers that meet our criteria. By successful replication, we mean that we can replicate the point estimate of the chosen specification up to the second decimal point. Figure A1 shows the number of replicable and non-replicable papers by year.
Figure A1: Replicability by Year
Note: The above figure shows the number of papers that meet our criteria. The grey and black bars represent the number of replicable papers and the number of papers that cannot be replicated.
A.4 Implementation Details
This section provides an elaboration of our reanalysis procedures documented in replication markdown files. For each paper, the reanalysis process encompasses five parts: (1) fundamental summary and visualization, (2) point estimates, (3) dynamic treatment effects, (4) diagnostic tests, and (5) sensitivity analyses.
A.4.1 Summary of Context and Visualization
Summary Table.
We meticulously document various aspects of each paper, including the outcome variable, treatment variable, unit and time indicators, covariates, treatment patterns, and fixed effects utilized.
Researchers primarily use two ways to motivate the TWFE model: they equate estimating a TWFE as employing a “difference-in-differences” (DID) strategy or they exploit “within”-unit (group) variations using a TWFE model.
We categorize the treatment pattern into three types as follows: (1) “Classic:” All treated units receive the treatment simultaneously, resembling a conventional difference-in-differences design; (2) “Staggered:” The treatment kicks in at different time points for different units and never reversal; and (3) “General:” The treatment can have reversals.
Visualizing Treatment Status:
We visualize the treatment status using the package panelView (Mou, Liu and Xu, 2023). The treated observations are visually represented by deep blue, whereas observations under control status are indicated by a lighter shade of blue. Additionally, we rearrange the units based on the chronological order of their initial exposure to the treatment.
Visualizing the Outcome:
Using the package panelView, we depict the trajectory of the outcome variable within the study’s time window for each individual unit. Control units are visually distinguished in gray, whereas treated units are represented by blue. Additionally, for studies with staggered adoption designs, we also plot the average outcome trajectory for each cohort.
A.4.2 Point Estimates
Original (Reported) Results:
We employ the fixest package (Berge, Krantz and McDermott, 2023) to run a fixed effect regression incorporating the treatment indicator, covariates, and fixed effects specified in the original specification. We print the raw regression output in the markdown file.
Replicated Results:
The replicated estimate are identical to the reported estimates except for Hall and Yoder (2022). Due to the large scale of the data from Hall and Yoder (2022), we base our analysis on a 1% sub-sample of the original data for computational efficiency. As a result, our estimate deviates slightly from the reported one. We present the unit-level clustered standard error and the standard error obtained from a 200-round clustered bootstrap, along with their corresponding confidence intervals.
Goodman-Bacon Decomposition:
For the analysis involving a staggered treatment pattern and no additional fixed effects apart from the unit and time fixed effects, we employ the Goodman-Bacon decomposition (Goodman-Bacon, 2021). This decomposition allows us to break down the replicated estimate into a weighted average of all possible $2\times 2$ DID estimates across different cohorts. Since the bacondecomp package is designed for balanced panel data, the weighted DID estimates may not perfectly align with the replicated estimate. Nonetheless, this approach serves as a valuable diagnostic tool to assess the potential impact of ‘invalid’ comparisons on the point estimate.
TWFE:
While maintaining the same regression specification as the replicated estimates, we modify the analysis by excluding always-treated units from the sample.
FEct:
We employ the fect package to implement the imputation methods. When the levels of fixed effects in the original specifications are higher than unit or time levels, we use the “cfe” method provided by the package for the imputation procedure. This is applied to three studies, including Bischof and Wagner (2019); Dewan, Meriläinen and Tukiainen (2020) and Schuit and Rogowski (2017). To obtain uncertainty estimates, we employ a 200-round cluster bootstrap approach. The program automatically excludes all always-treated observations.
Lagged Dependent Variables (LDVs):
In the estimation process using TWFE and FEct, we include the lagged-one-period outcome variable as a covariate. For studies with only two periods, we omit this step.
Unit-specific Linear Time Trends (ULT):
During the estimation process using TWFE and FEct, we introduce an additional fixed effect in the form of a unit-specific linear trend. If the original design already incorporates such trends, we omit this step to avoid redundancy.
Other HTE-Robust Estimators:
For all replications with no additional fixed effects apart from unit and time fixed effects, we implement the PanelMatch estimator (Imai, Kim and Wang, 2023). For analyses involving a staggered treatment pattern, we also implement multiple HTE-robust estimators, including the iw estimator (Sun and Abraham, 2021), the csdid estimator (Callaway and Sant’Anna, 2021). We also implement the stacked DID estimator for these staggered cases (Cengiz et al., 2019). We report the unit-level clustered standard error and the standard error derived from a 200-round clustered bootstrap, along with their corresponding confidence intervals. All always-treated units are always dropped automatically.
Stacked DID:
We adopt the methodology described in Bleiberg (2021) to implement the stacked DID estimator. We first construct a cohort-specific dataset for each ever-treated cohort, which includes the respective cohort and all never-treated units. Subsequently, these cohort-specific datasets are stacked to compute an average effect across all cohorts using a fixed effect regression model incorporating the treatment indicator, covariates, stack-unit interaction and stack-year interaction fixed effects.
iw Estimator:
We employ the sunab() command available in the fixest package to implement the iw estimator. To obtain the total average treatment effect, we set the att option to TRUE. We maintain the default values for other options.
csdid Estimator:
We utilize the did package to implement the csdid estimator (Sant’Anna and Callaway, 2021). We specifically set the est_method option to “reg,” employing only the outcome model rather than the double-robust model to estimate the ATT. Additionally, to compare the point estimate under different comparison group scenarios, we set the control_group option to both “notyettreated” and “nevertreated.”
PanelMatch:
For the analysis without additional fixed effects apart from the unit and time fixed effects, we use the PanelMatch package to implement the eponymous estimator. To determine the number of treatment history periods on which to matched, we set the lag option to the maximum value that does not result in an error from the command. By setting the match.missing option to TRUE, units are also matched based on the pattern of missingness in their treatment histories. Additionally, we set the covs.formula option to NULL and the refinement.method option to “none” to ensure equal weighting of control units within each matched set sharing the same treatment histories. The confidence interval is obtained from the built-in bootstrap method.
Balanced FEct:
By setting the option balance.period in fect, we are able to estimate the ATT for a specific subset of the sample using the imputation method. This subset comprises units with certain non-missing pretreatment and posttreatment periods. The lag and lead parameters employed here are kept consistent with the PanelMatch command.
A.4.3 Dynamic Treatment Effects & Event Study Plots
In the estimation of dynamic treatment effects, we designate the relative period 00 as the last pretreatment period and the relative period $1$ as the first posttreatment period. The rule of indexing the relative periods when treatment has reversals is clarified in Liu, Wang and Xu (2022). To implement this indexing rule, the get.cohort() command available in the paneltools package can be utilized. We use the command esplot() from the same package to visualize the DTE. All always-treated units are dropped in the estimation of DTE.
TWFE (No Reversals):
When the treatment doesn’t have reversals, we include a series of interaction terms between a dummy that indicates whether a unit is a treated unit and each lead (lag) indicator relative to the treatment in a fixed effect regression that incorporates the same fixed effects as specified in the original paper. We set the reference period as the last pretreatment period and obtain the confidence interval using clustered standard error and a 200-round clustered bootstrap.
TWFE (With Reversals):
When the treatment has reversals, we first use the get.cohort() command to determine the relative periods in relation to the treatment. We then generate the treated unit indicator “treat” using the following steps: First, for observations of units that have never been treated, we assign “treat” a value of 00. Second, for ever-treated units with treatment reversals, we assign “treat” a value of $1$ only for observations prior to the last treatment exit. For instance, if a unit’s treatment status sequence is $0,0,0,1,1,0,0$, we set “treat” to $1$ for the first five observations. Third, if certain units are already under the treated status in the initial period of the data, we exclude observations prior to their first treatment exit since their relative periods to treatment are uncertain. We then proceed to interact the binary variable “treat” with each lead (lag) indicator relative to the treatment and run the fixed-effect regression as previously described. We designate the last pretreatment period as the reference period and obtain the confidence interval using clustered standard error and a 200-round clustered bootstrap.
FEct:
The procedure here is the same as the one described in the previous section on point estimates.
For the analysis involving a staggered treatment pattern, we also implement the aforementioned HTE-robust estimators to estimate the DTE.
Stacked DID:
We create the cohort-specific datasets and the stacked data the same as we do in the point estimate section. The difference is that the fixed effect regression includes interaction terms between a dummy that indicates whether a unit is ever-treated and each lead (lag) indicator along with stack-unit interaction and stack-year interaction fixed effects. We designate the last pretreatment period as the reference period and obtain the confidence interval using clustered standard error and a 200-round clustered bootstrap.
iw Estimator:
We employ the same command as stated in the point estimate section, with the exception of setting the att option to FALSE. We aggregate the treatment effects to the relative period level without binning any periods. The reference period is still the last pretreatment period.
csdid Estimator:
We utilize a similar command to the one outlined in the point estimate section, with a few differences. Firstly, we specify the type option as “dynamic” to aggregate the treatment effects on the relative period level. Secondly, we set the cband option to FALSE to obtain period-wise confidence intervals. Lastly, we assign the base_period option as “universal” to establish the last pretreatment period as the base period.
PanelMatch:
We employ the same command as stated in the point estimate section, with the exception of setting placebo.test = TRUE to obtain the pseudo treatment effects for the pretreatment periods.
Balanced FEct:
The same as the point estimate.
A.4.4 Diagnostic Tests
For studies with sufficient numbers of pretreatment periods ($>3$), we employ the $F$-test, equivalence test, and placebo test to evaluate the validity of the parallel trend assumption (PTA). In the case of a study with treatment reversals and an adequate number of periods following the treatment exit ($>3$), we conduct a test to assess the absence of carryover effects. All these tests are based on estimations obtained from FEct, and information regarding the specific details of these tests can be found in Liu, Wang and Xu (2022). We report the $p$-values of these tests in the test results table.
$F$-Test:
We use an $F$-test to examine the presence of a pretrend. “Residuals” are defined as the differences between $Y(0)$ and $\hat{Y}(0)$. The null hypothesis posits that the residual averages in each pretreatment period are (jointly) equal to zero. This test encompasses all pretreatment periods wherein the count of ever-treated units exceeds 0.3 times the total number of treated units. A small $p$-value, leading to the rejection of the $F$-test, indicates the potential failure of the PTA.
Equivalence Test:
The equivalence test evaluates whether the 90% confidence intervals (corresponding to a 5% significance level) for the residuals in the pretreatment periods surpass a predetermined range, known as the equivalence range. The null hypothesis posits that the residual exceeds this specified range for each pretreatment period, thus a smaller $p$-value from the equivalence test indicates a better fit in pretreatment periods. We perform the equivalence test using two different pre-specified ranges. The first criterion uses the default range, set at $0.36\sigma_{\varepsilon}$, where $0.36\sigma_{\varepsilon}$ represents the standard deviation of the outcome variable partialling out the two-way fixed effects. The second criterion uses the estimated average treatment effect obtained through FEct. The $p$-values using the second criterion are reported in the main text.
Placebo Test:
We use the placebo test feature available in the fect package. We set the placebo.period option as c(-1,0), thereby excluding observations in the last two pretreatment periods during model fitting. We then examine whether the estimated ATT in these “placebo periods” significantly deviates from zero. The null hypothesis assumes that the average pseudo-treatment effects within this specified range are equal to zero. A small $p$-value, leading to the rejection of the placebo test, points to the potential failure of the PTA.
(No) Carryover Effects Test:
We utilize the (no) carryover test feature offered by the fect package. By specifying the carryover.period option as c(1,2), we exclude observations in the first two periods after exiting the treated status during model fitting. We then assess whether the estimated ATT in these periods significantly deviates from zero. The null hypothesis assumes that the average pseudo-treatment effect within this specified range is equal to zero. A small $p$-value, leading to the rejection of the carryover effects test, points to the potential failure of the assumption of no carryover effects.
A.4.5 Robust Confidence Set & Sensitivity Analysis
For studies with statistically significant ATT estimates given by FEct, we implement the sensitivity analysis developed by Rambachan and Roth (2023) to evaluate the robustness of the significant results under different extents of violation to PTA. Suppose the dynamic treatment effects $\beta$ in pretreatment placebo and posttreatment periods can be expressed as
$\beta=\underbrace{\left(\begin{array}[]{c}0\\ \tau_{\text{post}}\end{array}\right)}_{=:\tau\text{ treatment effects}}+\underbrace{\left(\begin{array}[]{c}\delta_{\text{placebo}}\\ \delta_{\text{post}}\end{array}\right)}_{=:\delta\text{ difference in trends}}$
where $\tau_{post}$ represents the treatment effects of interest and $\delta$ represents the difference in trends between the treated and comparison groups that would have occurred absent treatment. Following the practice in Rambachan and Roth (2023), we allow the existence of difference in trends ($\left(\begin{array}[]{c}\delta_{\text{placebo}}\\ \delta_{\text{post}}\end{array}\right)\neq 0$) but impose the restriction that the posttreatment violation of parallel trends between consecutive periods is no more than $\bar{M}$ times the maximum difference in trends between consecutive placebo periods, i.e,
$\delta\in\Delta^{RM}(\bar{M})=\left\{\delta:\forall t\geq 0,\left|\delta_{t+1}-\delta_{t}\right|\leq\bar{M}\cdot\max_{s\in\{-2,-1\}}\left|\delta_{s+1}-\delta_{s}\right|\right\},$
We use pretreatment placebo periods as the benchmark for difference in trends because placebo estimates are estimated by imputation, which are in the same way as posttreatment ATT estimates. With this restriction on $\delta$, we can use our estimated dynamic treatment effects $\hat{\beta}$ and observed difference in trends in placebo periods $\hat{\delta}_{placebo}=\hat{\beta}_{placebo}$ to construct the robust confidence intervals for the ATT and each posttreatment period as follows.
Point Estimates:
The ATT of interest can be expressed as $\theta=l_{att}^{\prime}\delta_{post}$, where $l_{att}=[\frac{n_{1}}{\sum_{t=1}^{T_{post}}n_{t}},\cdots,\frac{n_{T_{post}}}{\sum_{t=1}^{T_{post}}n_{t}}]^{\prime}$ and $n_{t}$ represents the number of observations at the posttreatment period $t$. By the lemma 2.1 in Rambachan and Roth (2023), define the set of values for $\theta$ that are consistent with a given value of $\beta$ under the restriction $\delta\in\Delta^{RM}(\bar{M})$ as $S(\beta,\Delta^{RM})=[\theta^{lb}(\beta,\Delta^{RM}),\theta^{ub}(\beta,\Delta^{RM})]$ where
$\theta^{lb}(\beta,\Delta^{RM}(\bar{M})):=l_{att}^{\prime}\beta_{\text{post}}-\left(\max_{\delta}l_{att}^{\prime}\delta_{\text{post}},\text{ s.t. }\delta\in\Delta^{RM}(\bar{M}),\delta_{\text{placebo}}=\beta_{\text{placebo}}\right)$
$\theta^{ub}(\beta,\Delta^{RM}(\bar{M})):=l_{att}^{\prime}\beta_{\text{post}}-\left(\min_{\delta}l_{att}^{\prime}\delta_{\text{post}},\text{ s.t. }\delta\in\Delta^{RM}(\bar{M}),\delta_{\text{placebo}}=\beta_{\text{placebo}}\right)$
Under a finite sample normal approximation for the estimated dynamic treatment effects $\hat{\beta}$, i.e., $\hat{\beta}\approx N(\tau+\delta,\Sigma_{n})$, Rambachan and Roth (2023) defines the confidence sets of $\theta$, $\mathcal{C}_{n}\left(\hat{\beta}_{n},\Sigma_{n}\right)$ by
$\inf_{\delta\in\Delta^{RM}(\bar{M}),\tau}\inf_{\theta\in\mathcal{S}(\delta+\tau,\Delta^{RM}(\bar{M}))}\mathbb{P}_{\hat{\beta}_{n}\sim\mathcal{N}\left(\delta+\tau,\Sigma_{n}\right)}\left(\theta\in\mathcal{C}_{n}\left(\hat{\beta}_{n},\Sigma_{n}\right)\right)\geq 1-\alpha$
They also show that under a large class of distributions $\mathcal{P}$ such that $\delta_{P}\in\Delta^{RM}(\bar{M})$ for all $P\in\mathcal{P}$, one can replace $\Sigma_{n}$ with a consistent estimate of $\hat{\Sigma}_{n}$ and have
$\liminf_{n\rightarrow\infty}\inf_{P\in\mathcal{P}}\inf_{\theta\in\mathcal{S}\left(\delta_{P}+\tau_{P},\Delta^{RM}(\bar{M})\right)}\mathbb{P}_{P}\left(\theta\in\mathcal{C}_{n}\left(\hat{\beta}_{n},\hat{\Sigma}_{n}\right)\right)\geq 1-\alpha$
We use the createSensitivityResults_relativeMagnitudes function in the Honestdid package to calculate $\mathcal{C}_{n}(\hat{\beta}_{n},\hat{\Sigma}_{n})$ for varying levels of $\bar{M}$. The estimated DTE $\hat{\beta}_{n}$ and the variance-covariance matrix $\hat{\Sigma}_{n}$ are computed using the fect function. To assess the robustness of the significant estimated ATT provided by FEct, we determine the threshold value of $\bar{M}$ where the corresponding confidence interval $\mathcal{C}_{n}(\hat{\beta}_{n},\hat{\Sigma}_{n})$ just includes 0.
Dynamic Treatment Effects:
We further scrutinize the robustness of the estimated DTE by computing the confidence interval at $\bar{M}=0.50$ for each posttreatment period. This analysis follows the same methodology as for the ATT, with the distinction lying solely in the definition of $l$.
A.5 More Replication Results
A.5.1 Reported vs TWFE and FEct Estimates
To assess how the imputation estimator differs from TWFE, we depict the ratio of the FEct estimates to the TWFE estimates, after excluding the always-treated units (ensuring identical sample sets), using solid circles in Figure A2. We also juxtapose the cited TWFE estimates with those omitting the always-treated units, represented by hollow circles. The red circles denote studies where the FEct estimates are not statistically significant at the 5% level. Notably, both the mean and median of these ratios are close to one. This finding suggests that, although there are noticeable differences in individual cases, TWFE does not systematically under- or over-estimates the ATT. Figure A2 also shows that the presence of always-treated units is not the primary driver of these differences. When these units are excluded, the TWFE estimates align closely with the reported estimates.
Figure A2: TWFE vs. Imputation Method: Estimates
Note: In the above figure, solid circles represent the ratios of the estimates from the imputation method (FEct) to TWFE coefficients with always-treated units removed; hollow circles represent the ratios of reported TWFE coefficients to TWFE coefficients with always-treated units removed. Statistically insignificant FEct estimates at the 5% level are painted in red.
The figures below show that (1) the presence of always-treated units does not significantly change the TWFE estimates; and (2) the application of FEct, an HTE-robust estimator, modestly changes the point estimates, resulting in a relatively large proportion of studies becoming statistical insignificant.
Figure A3: Reported, TWFE, and FEct Estimates: Full Sample
(a) Reported vs TWFE (alawys treated removed)
(b) TWFE (always treated removed) vs FEct
Note: The above figures compare reported coefficients, coefficients from a TWFE model with the authors’ preferred specification based on samples in which the always treated units are removed, and ATT estimates from FEct. The same reported SE estimate normalizes all three estimates in each application. Red triangles represent cases whose estimates on the y-axes are statistically insignificant at the 5% level
A.5.2 Inferential Methods
[…]